# Towards Energy-Aware Federated Learning via MARL: A Dual-Selection Approach for Model and Client

### v1

@@ Abstract.
Although Federated Learning (FL) is promising in knowledge sharing for heterogeneous Artificial Intelligence of Thing (AIoT) devices, their training performance and energy efficacy are severely restricted in practical battery-driven scenarios due to the “wooden barrel effect” caused by the mismatch between homogeneous model paradigms and heterogeneous device capability. As a result, due to various kinds of differences among devices, it is hard for existing FL methods to conduct training effectively in energy-constrained scenarios, such as battery constraints of devices. To tackle the above issues, we propose an energy-aware FL framework named DR-FL, which considers the energy constraints in both clients and heterogeneous deep learning models to enable energy-efficient FL. Unlike Vanilla FL, DR-FL adopts our proposed Muti-Agents Reinforcement Learning (MARL)-based dual-selection method, which allows participated devices to make contributions to the global model effectively and adaptively based on their computing capabilities and energy capacities in a MARL-based manner. Experiments on various well-known datasets show that DR-FL can not only maximize knowledge sharing among heterogeneous models under the energy constraint of large-scale AIoT systems but also improve the model performance of each involved heterogeneous device.
@@ 1. Introduction
As a promising collaborative machine learning paradigm, FL allows local DL model training among various devices without compromising their local data privacy. Instead of sharing local sensitive data among devices, FL only needs to send gradients or weights of local device models to a cloud server for knowledge aggregation, thus enhancing both the training and inference capability of local models. Although FL is promising in knowledge sharing, it faces the problems of both large-scale deployment and quick adaption to dynamic environments, where local models are required to be frequently trained to accommodate an ever-changing world. In practice, such problems are hard to be solved, since classic Federated Averaging (e.g., FedAvg) methods require that all devices should have homogeneous local models with the same architecture.
According to the well-known “wooden barrel effect” caused by homogeneous assumption as shown in Figure 1, the energy consumption waste in Vanilla FL is usually due to the following two reasons, i.e., the mismatch between computing power and homogeneous model, and the mismatch between power consumption and homogeneous model. The former uses device energy for waiting time, while the latter uses device energy for useless training time (only enough power to support training but not support communication). Thus, such a homogeneous model assumption strongly limits the overall energy efficiency of the entire FL system. This is because energy usages in the entire system are mainly determined by how much power is used in the effective model learning other than waiting, which consumes energy to wait other than training or communication.
@@
Although various heterogeneous FL methods (e.g., HeteroFL (Diao et al., 2021), Scale-FL (Ilhan et al., 2023), PervasiveFL (Xia et al., 2022)) and energy-saving techniques (Li et al., 2020, 2019) have been investigated to address the above issue, most of them focus on either enabling effective knowledge sharing between heterogeneous models or reducing the energy consumption of devices. Based on the coarse-grained FedAvg operations, few of the existing FL methods can substantially address the above challenges to quickly adapt to new environments within an energy-constrained scenario. Inspired by the concepts of BranchyNet (Teerapittayanon et al., 2016) and multi-agent reinforcement learning (Zhang et al., 2021a), in this paper, we propose a novel FL framework named DR-FL, which takes both the layer-wise structure information of DL models and the remaining energy of each client into account to enable energy-efficient federated training. Unlike traditional FedAvg-based FL method that relies on homogeneous device models, DR-FL maintains a layer-wise global model on the cloud server, while each device only installs a subset layer-wise model according to its computing power and remaining battery. In this way, all the heterogeneous local models can effectively make contributions to the global model based on their computing capabilities and remaining energy in a MARL-based manner. Meanwhile, by adopting MARL, DR-FL can not only make the trade-off between training performance and energy consumption, thus ensuring energy-efficient FL training to accommodate various energy-constrained environments. This paper makes the following three major contributions:
  • We establish a novel, lightweight cloud-based FL framework named DR-FL, which can be easily implemented and enables various heterogeneous DNNs to share knowledge without compromising their data privacy in FL for heterogeneous devices by layer-wise model aggregation.
@@ 2. Related Work
Although FL is good at knowledge sharing without compromising the data privacy of devices in AIoT system design, due to the homogeneous assumption that all the involved devices should have local DL models with the same architecture, Vanilla FL methods inevitably suffer from the problems of low inference accuracy and invalid energy consumption, thus impeding the deployment of FL methods in large-scale AIoT system designs (Khan et al., 2020; Nguyen et al., 2021; Xia et al., 2022; Zhu et al., 2021), especially for non-IID scenarios.
To enable collaborative learning among heterogeneous device models, various solutions have been extensively studied, which can be primarily classified into two categories, i.e., subnetwork aggregation-based methods and knowledge distillation-based methods. The basic idea of subnetwork aggregation-based methods is to allow knowledge aggregation on top of subnetworks of local device models, which enables knowledge sharing among heterogeneous device models. For instance, Diao et al. (Diao et al., 2021) presented an effective heterogeneous FL framework named HeteroFL, which can train heterogeneous local models with varying computation complexities but still produce a single global inference model, assuming that device models are subnetworks of the global model. By integrating FL and width-adjustable slimmable neural networks, Yun et al. (Yun et al., 2023) proposed a novel learning framework named ScaleFL, which jointly utilizes superposition coding for global model aggregation and superposition training for updating local models. In (Xia et al., 2022), Xia et al. developed a novel framework named PervasiveFL, which utilizes a small uniform model (i.e., “modellet”) to enable heterogeneous FL. Although all the above heterogeneous FL methods are promising, most focus on improving inference performance. Few of them take the issues of real-time training and energy efficiency into account.
@@
To the best of our knowledge, DR-FL is the first attempt to investigate the dual selection by both layer-wise models and the participated clients based on MARL to enable fine-grained heterogeneous FL, where heterogeneous devices can adaptively and efficiently make contributions to the global model based on their computing capabilities and remaining energy. Compared with state-of-the-art heterogeneous FL methods, DR-FL can not only maximize the knowledge sharing among various heterogeneous models under energy constraints but also significantly improve both the model performance of each involved device and the energy efficacy of the entire FL system.
@@ 3.1. Federated Learning
where $\mathbb{W}_{t}^{n}$ and $\mathbb{W}_{t+1}^{n}$ represent the global models at round $t$ and round $t+1$ in the $n^{th}$ device, respectively. $\eta$ indicates the learning rate and $\nabla\mathbb{W}_{t}^{n}$ is the gradient obtained by the $n^{th}$ device model after the $t^{th}$ training round. To protect data privacy, at the end of each communication round, FL uploads the weight differences (i.e., model gradients) of each device instead of the newly updated models to the cloud for aggregation. After gathering the gradients from all the participating devices, the cloud updates the parameters of the shared-global model based on the Fedavg (McMahan et al., 2016b) algorithm, which is defined as follows:
@@
where $\frac{\sum_{n=1}^{N}\nabla\mathbb{W}_{t}^{n}}{N}$ denotes the average gradient of $N$ participating devices in communication round $t$, $\mathbb{W}_{t}$ and $\mathbb{W}_{t+1}$ represent the global models after $t^{th}$ and $t+1^{th}$ communication round, respectively, and $\mathbb{L}_{n}$ means the training data size of device $n$. Although Vanilla FL methods (e.g., FedAvg) perform remarkably in distributed machine learning, they cannot be directly applied to AIoT scenarios. This is because the heterogeneous AIoT devices will lead to different training speeds for Vanilla FL, resulting in additional energy waste, which is unacceptable for an energy-constrained system.
@@ 3.2. Multi-Agent Reinforcement Learning
In cooperative Multi-Agent Reinforcement Learning (MARL), a set of $N$ agents is trained to produce optimal actions that lead to maximum team rewards. Specifically, at each timestamp $t$, each agent $n$ (where $1\leq n\leq N$) observes its state $s_{t}^{n}$ and selects an action $a_{t}^{n}$ based on $s_{t}^{n}$. After all agents have completed their actions, the team receives a joint reward $r_{t}$ and transitions to the next state $s_{t+1}^{n}$. The goal is to maximize the total expected discounted reward $R=\sum_{t=1}^{T}\gamma r_{t}$ by selecting the optimal actions for each agent, where $\gamma\in[0,1]$ is the discount factor.
Recently, QMIX (Rashid et al., 2018) has emerged as a promising solution for jointly training agents in cooperative MARL. In QMIX, each agent $n$ employs a Deep Neural Network (DNN) to infer its actions. This DNN implements the $Q$-function $Q^{\theta}(s,a)=E[R_{t}|s_{t}^{n}=s,a_{t}^{n}=a]$, where $\theta$ represents the parameters of the DNN, and $R_{t}=\sum_{i=t}^{T}\gamma r_{i}$ is the total discounted team reward received at $t$. During MARL execution, each agent $n$ selects the action $a^{*}$ with the highest $Q$-value (i.e., $a^{*}=\arg\max_{a}Q^{\theta}(s_{n},a)$).
To train the QMIX, a replay buffer is employed to store transition tuples $(s_{t}^{n},a_{t}^{n},s_{t+1}^{n},r_{t})$ for each agent $n$. The joint $Q$-function, $Q_{\text{tot}}(\cdot)$, is represented as the element-wise summation of all individual $Q$-functions (i.e., $Q_{\text{tot}}(s_{t},a_{t})=\sum_{n}Q^{\theta}_{n}(s_{t}^{n},a_{t}^{n})$), where $s_{t}=\{s_{t}^{n}\}$ and $a_{t}=\{a_{t}^{n}\}$ are the states and actions collected from all agents $n\in N$ at timestamp $t$. The agent DNNs can be recursively trained by minimizing the loss $L=E_{s_{t},a_{t},r_{t},s_{t+1}}[y_{t}-Q_{\text{tot}}(s_{t},a_{t})]^{2}$, where $y_{t}=r_{t}+\gamma\sum_{n}\max_{a}Q^{\theta^{\prime}}_{n}(s_{t+1}^{n},a)$ and $\theta^{\prime}$ represents the parameters of the target network, which are periodically copied from $\theta$ during the training phase.
@@ 4.1. Problem Formulation
Assuming that an energy-constrained FL system contains a cloud server and $N$ heterogeneous AIoT devices, which can be represented as $D=\left\{D_{1},...,D_{n},...,D_{N}\right\}$. All these heterogeneous AIoT devices can be classified into three categories according to their computing capability, i.e., small, middle and large, where the small, middle and large mean the level of device computing resources and storage resources. In this paper, the performance of the entire FL system is significantly influenced by three key factors: running time, energy consumption, and model accuracy. Running time determines the training efficiency of the FL system in a real scenario. Moreover, energy consumption is also a significant factor, particularly for AIoT devices powered by limited energy resources. Lastly, model accuracy ensures that the system produces reliable and valuable predictions. Therefore, to optimize the overall performance of the FL system, it is crucial to make a balance between three factors.
@@
Model Accuracy: In a heterogeneous scenario, how to effectively leverage the heterogeneity in heterogeneous models and devices to enhance the performance of aggregated models is an urgent issue that needs to be solved in FL. Furthermore, resource-constrained heterogeneous AIoT devices that participate in aggregation pose a considerable obstacle to the application of energy-constrained FL. Inspired by the work (Li et al., 2019) where the performance of model inference is affected by the number of successful aggregations for its device, we can deduce that the accuracy of heterogeneous models is proportional to the total number of aggregated models participating in each round. However, since devices consume energy every time they participate in each round of aggregation, how to reasonably select aggregation devices in an energy-constrained environment to improve model accuracy has become a major challenge in the design of an FL framework.
Optimization Objective: Taking energy information into account, an optimization model is proposed for energy-constrained FL. This model aims to balance three objectives: minimizing total running time $T_{all}$, and maximizing model accuracy $M_{acc}$ under total energy consumption $E_{all}$ constraint, which is defined as follows
(8) $\begin{matrix}\min T_{all},\ \quad\max M_{acc},\\ \text{s.t.}\ E_{all}\leq E,\end{matrix}$
where $E$ is the energy budget of an FL system.
@@ 4.2. Workflow of DR-FL
In DR-FL, heterogeneous AIoT devices and a cloud server cooperate to achieve high performance of various layer-wise models deployed on edge devices. Before training, all devices participating in DR-FL will initialize and install a layer-wise model, which is a subset layer of the global model in the cloud server. Then, the cloud server sends a part of the global model to AIoT devices for local training. At the end of local training, DR-FL performs layer-wise model aggregation on the cloud server. Note that hot-plug AIoT devices are permissible in DR-FL, where newly involved devices only inherit the parameters of the global model in the cloud server. Figure 2 shows the workflow of the DR-FL, which consists of five steps as follows.
Step 1 (Battery or Model Information Upload): During the initialization step of DR-FL, each device intending to participate in FL should upload its device information to the cloud, which includes the power, computing, and storage capabilities of devices and the overclocking potential of models. This collected information is used for energy-aware dual-selection for the layer-wise model and client in subsequent steps to optimize the entire system’s energy efficiency.

### v2

@@ Abstract.
Although Federated Learning (FL) is promising in knowledge sharing for heterogeneous Artificial Intelligence of Thing (AIoT) devices, their training performance and energy efficacy are severely restricted in practical battery-driven scenarios due to the “wooden barrel effect” caused by the mismatch between homogeneous model paradigms and heterogeneous device capability. As a result, due to various kinds of differences among devices, it is hard for existing FL methods to conduct training effectively in energy-constrained scenarios, such as battery constraints of devices. To tackle the above issues, we propose an energy-aware FL framework named DR-FL, which considers the energy constraints in both clients and heterogeneous deep learning models to enable energy-efficient FL. Unlike Vanilla FL, DR-FL adopts our proposed Muti-Agents Reinforcement Learning (MARL)-based dual-selection method, which allows participated devices to make contributions to the global model effectively and adaptively based on their computing capabilities and energy capacities in a MARL-based manner. Experiments conducted with various widely recognized datasets demonstrate that DR-FL has the capability to optimize the exchange of knowledge among diverse models in large-scale AIoT systems while adhering to energy limitations. Additionally, it improves the performance of each individual heterogeneous device’s model.
@@ 1. Introduction
As a promising collaborative machine learning paradigm, FL allows local DL model training among various devices without compromising their local data privacy. Instead of sharing local sensitive data among devices, FL only needs to send gradients or weights of local device models to a cloud server for knowledge aggregation, thus enhancing both the training and inference capability of local models. Although FL is promising in knowledge sharing, it faces the problems of both large-scale deployment and quick adaption to dynamic environments, where local models are required to be frequently trained to accommodate an ever-changing world. In practice, such problems are hard to be solved, since vanilla FL methods require that all devices should have homogeneous local models with the same architecture.
According to the well-known “wooden barrel effect” caused by homogeneous assumption as shown in Figure 1, the energy consumption waste in vanilla FL is usually due to the following two reasons, i.e., the mismatch between heterogeneous computing power and homogeneous model, and the mismatch between heterogeneous power consumption and homogeneous model. The former uses device energy for waiting time, while the latter uses device energy for useless training time (only enough power to support training but not support communication). Thus, such a homogeneous model assumption strongly limits the overall energy efficiency of the entire FL system. This is because energy usages in the entire system are mainly determined by how much power is used in the effective model learning other than waiting or useless training, which consumes energy to wait other than training or communication.
@@
Although various heterogeneous FL methods (e.g., HeteroFL (Diao et al., 2021), Scale-FL (Ilhan et al., 2023), PervasiveFL (Xia et al., 2022)) and energy-saving techniques (Li et al., 2020, 2019) have been investigated to address the above issue, most of them focus on either enabling effective knowledge sharing between heterogeneous models or reducing the energy consumption of devices. Based on the coarse-grained FedAvg operations, few of the existing FL methods can substantially address the above challenges to quickly adapt to new environments within an energy-constrained scenario. Inspired by the concepts of BranchyNet (Teerapittayanon et al., 2016) and multi-agent reinforcement learning (Zhang et al., 2021a), in this paper, we propose a novel FL framework named DR-FL, which takes both the layer-wise structure information of DL models and the remaining energy of each client into account to enable energy-efficient federated training. Unlike the traditional FL method that relies on homogeneous device models, DR-FL maintains a layer-wise global model on the cloud server, while each device only installs a subset layer-wise model according to its computing power and remaining battery. In this way, all the heterogeneous local models can effectively make contributions to the global model based on their computing capabilities and remaining energy in a MARL-based manner. Meanwhile, by adopting MARL, DR-FL can not only make the trade-off between training performance and energy consumption, thus ensuring energy-efficient FL training to accommodate various energy-constrained environments. This paper makes the following three major contributions:
  • We establish a novel lightweight cloud-based FL framework named DR-FL, which can be easily implemented and enables various heterogeneous DNNs to share knowledge without compromising their data privacy in FL for heterogeneous devices by layer-wise model aggregation.
@@ 2. Related Work
Although FL is good at knowledge sharing without compromising the data privacy of devices in AIoT system design, due to the homogeneous assumption that all the involved devices should have local DL models with the same architecture, vanilla FL methods inevitably suffer from the problems of low inference accuracy and invalid energy consumption, thus impeding the deployment of FL methods in large-scale AIoT system designs (Khan et al., 2020; Nguyen et al., 2021; Xia et al., 2022; Zhu et al., 2021), especially for non-IID scenarios with constrained energy limitation.
To facilitate collaborative learning among different device models, many works have thoroughly examined numerous solutions. These solutions may be broadly categorized into two types: subnetwork aggregation-based methods and knowledge distillation-based methods. Subnetwork aggregation-based methods aim to facilitate knowledge aggregation by aggregating subnetworks of local device models. This approach enables the sharing of knowledge among diverse device models. For instance, Diao et al. (Diao et al., 2021) presented an effective heterogeneous FL framework named HeteroFL, which can train heterogeneous local models with varying computation complexities but still produce a single global inference model, assuming that device models are subnetworks of the global model. By integrating FL and width-adjustable slimmable neural networks, Yun et al. (Yun et al., 2023) proposed a novel learning framework named ScaleFL, which jointly utilizes superposition coding for global model aggregation and training for updating local models. In (Xia et al., 2022), Xia et al. developed a novel framework named PervasiveFL, which utilizes a small uniform model (i.e., “modellet”) to enable heterogeneous FL in AIoT systems. Although all the above heterogeneous FL methods are promising, most of them focus on improving inference performance of local models. Few of them take the issues of real-time training and energy efficiency into account.
@@
To the best of our knowledge, DR-FL is the first attempt to investigate the dual selection by both layer-wise models and the participated clients based on MARL to enable fine-grained heterogeneous FL, where heterogeneous devices can adaptively and efficiently make contributions to the global model based on their computing capabilities and remaining energy. DR-FL surpasses other advanced heterogeneous FL approaches by optimizing information transfer among different models with limited energy resources and enhancing both the performance of individual devices and the energy efficiency of the entire FL system.
@@ 3.1. Federated Learning
where $\mathbb{W}_{t}^{n}$ and $\mathbb{W}_{t+1}^{n}$ represent the global models at round $t$ and round $t+1$ in the $n^{th}$ device, respectively. $\eta$ indicates the learning rate and $\nabla\mathbb{W}_{t}^{n}$ is the gradient obtained by the $n^{th}$ device model after the $t^{th}$ training round. To preserve the data privacy of local devices, at the end of each communication round, FL uploads each device’s weight differences (i.e., model gradients) instead of updated NN models to the cloud for aggregation. After gathering the gradients from all the participating devices, the cloud updates the parameters of the shared-global model based on the Fedavg (McMahan et al., 2016b) algorithm, which is defined as follows:
@@
where $\frac{\sum_{n=1}^{N}\nabla\mathbb{W}_{t}^{n}}{N}$ denotes the average gradient of $N$ participating devices in communication round $t$, $\mathbb{W}_{t}$ and $\mathbb{W}_{t+1}$ represent the global models after $t^{th}$ and $t+1^{th}$ communication round, respectively, and $\mathbb{L}_{n}$ means the training data size of device $n$. Although vanilla FL methods (e.g., FedAvg) perform remarkably in distributed machine learning, they cannot be directly applied to AIoT scenarios. This is because the heterogeneous AIoT devices will lead to different training speeds for vanilla FL, resulting in additional energy waste, which is unacceptable for an efficient energy-constrained system.
@@ 3.2. Multi-Agent Reinforcement Learning
Cooperative Multi-Agent Reinforcement Learning (MARL) involves training a group of $N$ agents to generate optimal actions that result in the highest possible team rewards. At each timestamp $t$, each agent $n$ (where $1\leq n\leq N$) observes its state $s_{t}^{n}$ and chooses an action $a_{t}^{n}$ based on $s_{t}^{n}$. Once all agents have finished their actions, the team is given a collective reward $r_{t}$ and moves on to the next state $s_{t+1}^{n}$. The objective is to optimize the overall predicted discounted reward $R=\sum_{t=1}^{T}\gamma r_{t}$ by choosing the best behaviours for each agent. Here, $\gamma\in[0,1]$ represents the discount factor for reward.
Recently, QMIX (Rashid et al., 2018) has emerged as a promising solution for jointly training agents in cooperative MARL. In QMIX, each agent $n$ employs a Deep Neural Network (DNN) to infer its actions. This DNN implements the $Q$-function $Q^{\theta}(s,a)=E[R_{t}|s_{t}^{n}=s,a_{t}^{n}=a]$, where $\theta$ represents the parameters of the DNN, and $R_{t}=\sum_{i=t}^{T}\gamma r_{i}$ is the total discounted team reward received at $t$. During MARL execution, each agent $n$ selects the action $a^{*}$ with the highest $Q$-value (i.e., $a^{*}=\arg\max_{a}Q^{\theta}(s_{n},a)$). To train the QMIX, a replay buffer is employed to store transition tuples $(s_{t}^{n},a_{t}^{n},s_{t+1}^{n},r_{t})$ for each agent $n$. The joint $Q$-function, $Q_{\text{tot}}(\cdot)$, is represented as the element-wise summation of all individual $Q$-functions (i.e., $Q_{\text{tot}}(s_{t},a_{t})=\sum_{n}Q^{\theta}_{n}(s_{t}^{n},a_{t}^{n})$), where $s_{t}=\{s_{t}^{n}\}$ and $a_{t}=\{a_{t}^{n}\}$ are the states and actions collected from all agents $n\in N$ at timestamp $t$. The agent DNNs can be recursively trained by minimizing the loss $L=E_{s_{t},a_{t},r_{t},s_{t+1}}[y_{t}-Q_{\text{tot}}(s_{t},a_{t})]^{2}$, where $y_{t}=r_{t}+\gamma\sum_{n}\max_{a}Q^{\theta^{\prime}}_{n}(s_{t+1}^{n},a)$ and $\theta^{\prime}$ represents the parameters of the target network, which are periodically copied from $\theta$ during the training phase.
@@ 4.1. Problem Formulation
Given an energy-limited federated learning (FL) system consisting of a cloud server and $N$ diverse AIoT devices, denoted as $D=\left\{D_{1},...,D_{n},...,D_{N}\right\}$. These diverse AIoT devices can be categorized into three separate categories based on their computing capability: small, medium, and large. The terms small, medium, and large refer to the level of computational and storage capabilities available on the devices. Three crucial aspects, namely running time, energy consumption, and model correctness, substantially impact the overall performance of the FL system discussed in this paper. The running time of the FL system directly impacts the training efficiency in a real-world scenario. Furthermore, energy consumption is a crucial aspect, especially for AIoT devices that rely on limited energy resources. Finally, the accuracy of the model guarantees that the system generates reliable and valuable predictions. Hence, to enhance the FL system’s overall efficiency, achieving a harmonious equilibrium among three key components is imperative.
@@
Model Accuracy: The appropriate utilization of heterogeneity in heterogeneous models and devices to improve the performance of aggregated models is an urgent issue that must be addressed in the field of Federated Learning (FL). In addition, the use of energy-limited federated learning is hindered by the presence of resource-constrained heterogeneous AIoT devices that are involved in data aggregation. Based on the findings of the study referenced as (Li et al., 2019), it can be inferred that the accuracy of heterogeneous models is directly related to the number of successful aggregations for each device. In other words, the more aggregated models participating in each round, the higher the accuracy of the model inference. Nevertheless, the issue in designing a Federated Learning (FL) framework lies in the selection of aggregation devices in an energy-constrained environment to enhance model accuracy, considering that devices waste energy during each cycle of aggregation.
Optimization Objective: Taking energy information into account, an optimization model is proposed for energy-constrained FL. This model seeks to achieve a compromise between three objectives: minimizing the overall running time $T_{all}$, maximizing the model accuracy $M_{acc}$ while adhering to a constraint on total energy consumption $E_{all}$. The constraint is defined as follows:
(8) $\begin{matrix}\min T_{all},\ \quad\max M_{acc},\\ \text{s.t.}\ E_{all}\leq E.\end{matrix}$
Here, $E$ represents the energy allocation of a FL system.
@@ 4.2. Workflow of DR-FL
DR-FL involves the collaboration of diverse AIoT devices and a cloud server to optimize the performance of different layer-wise models implemented on edge devices. Prior to training, all devices involved in DR-FL will initialize and install a layer-wise model, which is a subset layer of the global model stored on the cloud server. Subsequently, the cloud server transmits a segment of the comprehensive model to AIoT devices for localized training. After completing the training process on the local device, DR-FL carries out layer-wise model aggregation on the cloud server. It should be noted that hot-plug AIoT devices are allowed with DR-FL. These newly connected devices only receive the parameters of the global model from the cloud server. Figure 2 depicts the process of the DR-FL, which comprises five serial steps.
Step 1 (Battery or Model Information Upload): During the initialization step of DR-FL, each device intending to participate in FL should upload its device information to the cloud, which includes the power, computing, and storage capabilities of devices and the overclocking potential of models. In subsequent steps, this information is used for energy-aware dual-selection for the layer-wise model and client to optimize the entire system’s energy efficiency.